# Technology, Institution, and Regional Growth: Evidence from Mineral Mining Industry in Industrializing Japan

### v12

@@ 4.3 Dependent Variables
The primary dependent variable, population, is obtained by digitizing the municipal-level statistics documented in the 1930 and 1935 Population Censuses (Statistics Bureau of the Cabinet 1935b, 1939).
@@
To investigate the natural change mechanism behind the local population growth due to mining, I also digitized the Municipal Vital Statistics of 1930 and 1935 (Statistics Bureau of the Cabinet 1932c, 1938), which document the total number of marriages, live births, and deaths in all municipalities in the survey years. I then consider four dependent variables: crude marriage rate, crude birth rate, marital fertility rate, and gender-specific death rate. [^26] I followed the conventional definitions of these variables. The crude marriage and birth rates are the number of marriages and live births per $1,000$ people, respectively. Marital fertility is defined as the number of live births per $1,000$ married females. Although the number of married females in each age is unavailable, the number of married females is similar to the number of households in each municipality, meaning that the marital fertility definition herein is useful in representing the couple’s fertility decision. In other words, the results from this marital fertility definition shall not be sensitive to the existence of the elder married women in the households. Although marital fertility is considered to capture the couple’s fertility decision better, the results can be materially similar between crude and marital fertilities because out-of-wedlock births have been historically rare in Japan. The crude death rate is the number of deaths per $1,000$ people. These are used to test the fertility responses of the miners and the changes in mortality due to mining activities and regulations. To further delve into the occupational hazard channel, I consider the infant mortality rate and fetal death rate. [^27] The infant mortality rate is the number of infant deaths per $1,000$ live births. The number of censored observations in infant mortality is negligible. The fetal death rate is defined as the number of fetal deaths per $1,000$ births. The censored observations in the fetal death rates are less than $11$% in both census years. The estimates from the Tobit estimator (Tobin 1958) are materially similar to those from the OLSE (not reported). This confirms the evidence that censoring does not matter in my empirical setting. These rates allow for testing the impacts of the occupational hazards for female miners on the mortality selection before birth under the biological framework. I also consider the mortality rate of children aged 1–4 to assess the potential impacts of air pollution. [^28] The child mortality rate is defined as the number of deaths of children aged 1–4 in 1933 (or 1938) per $1,000$ children aged 1–5 in 1930. Note that the number of children aged 1–5 from 1930 is used as it is only documented in the 1930 population census. Given that child deaths, especially at age 5, were rarer than infant deaths at that time, the mismatch in the age bins between numerator and denominator should not be a critical issue. The difference in the survey points does add noise to the estimation of the standard error. However, the coefficient estimate is close enough to zero and thus, should not be a practical issue in this setting (Table 5). To obtain both variables, I digitized the 1933 and 1938 reports documenting municipal-level statistics on births and infant deaths published by the Aiikukai and Social Welfare Bureau.
@@
Comprehensive statistics on migrants were unavailable in the prewar period. As described earlier, however, once the contribution of natural change to population growth is known, the sign of social change in the geometric population growth model is uniquely determined. Considering this, I rather assess to what extent the influx of single male miners disturbed the gender balance and average household size and how it changed after the regulation was enforced. The data on the sex ratio and average household size are obtained from the official reports of the 1930 and 1935 Population Censuses. [^29] I followed the conventional definitions on these variables. The sex ratio is the number of males divided by the number of females. The average household size is the number of people in households divided by the number of households. Overall, these variables are expected to represent the characteristics of local demographics (Angrist 2002; Abramitzky et al. 2011).
@@
Importantly, the census documents the total number of people in all municipalities in the census years. Similarly, the vital statistics are corrected under a national registration system that covers all the births and deaths in a certain year. [^30] The potential imprecision of birth data in prewar Japan was mostly eliminated by the 1920s (Drixler 2016). This comprehensiveness helps avoid sample selection issues in statistical inference. Online Appendix Table C.4 presents the summary statistics on these variables by treatment status and census year. The mean differences are statistically significant for most variables, implying that systematic demographic changes may have occurred around the mining extraction area.
@@ 4.4 Control Variables
As coal is a bulky mineral, railways were the primary transportation of the extracted coal. Firms may have preferred setting extraction points closer to railway stations to reduce transportation costs (Sumiya 1968, pp. 440–441). Further, the local population size could be larger if one moves closer to stations. Therefore, controlling for the distance between each municipality and the nearest-neighboring station is preferable to deal with the potential endogeneity issue. I used the location information of all stations obtained from the official dataset provided by the Ministry of Land, Infrastructure, Transport and Tourism to compute the nearest-neighbor distance to stations in 1931 and 1936.
Although railways were primarily used to transport coal because coal mines were usually inland, marine transportation was also used for secondary logistics (Online Appendix Figure 2). Hence, similar to railway transportation, I include the distance to the nearest neighboring seaport as a variable to control for accessibility to marine transportation.
I also consider the accessibility to rivers. Not all mining areas have large rivers suitable for water transportation, meaning that rivers did not dominate the locations of coal mines. However, several mines used river transport as primary logistics until the development of railways before the late 1920s (Chiba and Yamada 1964). This implies that the spatial distribution of the river might have influenced the location of mines in such mining areas. To control for this possibility, I included the distance to the nearest neighboring river as a control variable.
Finally, I take the potential influence of the elevation into account. The location of the mine might have been correlated with the elevation of municipalities because terrain may relate to the locational fundamentals such as the climatic conditions and fixed cost of transportation. I used the grid cell elevation data provided by the Ministry of Land, Infrastructure, Transport, and Tourism to systematically calculate the average elevation in each municipality.
As explained below, including these control variables has little influence on the estimates after conditioning on the city and town fixed effects. To be conservative, however, I prefer to include all the controls to ensure the orthogonality condition of the instrumental variable. Online Appendices C.4.1C.4.4 summarizes the details of the data on railways, seaports, rivers, and elevation. Online Appendix Table C.5 presents the summary statistics on these control variables. [^31] Online Appendix Figures C.5C.6, and C.7 illustrate the locations of the railway stations, seaports, and rivers in the Japanese archipelago, respectively. Online Appendix Figure C.8 shows the spatial distribution of the elevation. All the controls are log-transformed to better model the skewness of the distances and elevations in nature, albeit this transformation does not influence the results.
@@ 5.1 Estimation Strategy
I leverage the random nature of mineral deposits to identify the effects of mines. The linear regression model is as follows:
@@
where $i$ indexes municipalities, $y$ is the outcome variable, MineDeposit is an indicator variable that equals one for municipalities within 5 km from a mine, $\mathbf{x}$ is a vector of control variables, and $e$ is a random error term. Because the placement of a mine is determined by a geological anomaly, MineDeposit is random in nature (Benshaul-Tolonen 2019, p.1568). However, the rest of the variation may be correlated with the local variation in infrastructure. For instance, if a mineral deposit is found between a village and a city, the mining firm may choose the city as its main mining point because the city is likely to have better infrastructure than the village. Then, MineDeposit can be positively correlated with the error term, leading to a positive (negative) omitted variable bias in the estimate of $\beta$ if the placement of the city is positively (negatively) correlated with the outcome variable.323232 For example, consider a simplified projection of $y$ on MineDeposit and City, $y=\alpha+\beta\text{{MineDeposit}}+\gamma\textit{City}+e$. When the municipality type (City) is unobserved, the linear projection coefficient can be written as $\beta^{*}=\beta+\Xi\gamma$, where $\Xi=(E[\text{{MineDeposit}}^{2}]^{-1})(E[\text{{MineDeposit}}\text{{City}})$. As explained, MineDeposit and City may be positively correlated such that $\Xi>0$. Thus, when $y$ is the population, conditional on mine deposits, it is reasonable to suppose that the town has a greater number of people than villages ($\gamma>0$). This result implies that $\beta^{*}=\beta+\Xi\gamma>\beta$. To deal with this systematic bias in the ordinary least squares, I included two indicator variables that equal one for cities and towns, respectively. As explained in Section 4.4, I further include the distances to the nearest neighboring railway station, seaport, and river and the elevation to explicitly control for transportation accessibility. [^33] The differences in the estimates from both specifications, including and excluding the control variables, indicate how much this omitted variable mechanism influences the results. Thus, I can partially assess the randomness of the primary exposure variable (MineDeposit). As I show later, the results from the specification including the city and town fixed effects are materially similar to those from the specifications, including the control variables, thereby supporting the randomness of the exposure variable. To be conservative, I prefer the specification including the control variables.
@@
A potential threat in the identification may be measurement errors in time-dimensional assignments. As explained in Section 4.1, mine data were surveyed in October 1931 (1936), whereas the Population Census was conducted in October 1930 (1935). For this lag concerning the censuses should not lead to a strong attenuation issue because the increments in the number of mines per year were considerably limited (Section 4.1). In turn, the data on fetal death and infant mortality were obtained from the 1933 (1938) Vital Statistics. The lags in the matching allow me to consider the exposure durations. For instance, fetus miscarriages in 1933 (1938) were in utero conceived in 1932 (1937), whereas their mothers should have been exposed to any shocks at the time of conception in 1932 (1937). The same argument can be applied to infant mortality: infants who passed away in 1933 (1938) were born at the beginning of 1932 (1937) at the earliest. Therefore, they should have been in utero in 1931 (1936). Consequently, the mining data that list the mines in October 1931 and 1936 may be reasonably matched with the health data of 1933 and 1938, respectively. However, one must be careful as there is still a one-year lag in matching with both the census and vital statistics datasets. This may lead to attenuation bias due to miss assignments in the time dimension, because the number of coal mines should be increased during that year. [^34] Note that this sort of measurement error never overstates the estimates but causes attenuation. For instance, if a few treated (untreated) municipalities were regarded as the untreated (treated) municipalities, the impacts of mines shall be discounted. See Online Appendix D.1 for a brief explanation of this mechanism.
@@
To deal with such potential issues, I consider the IV estimator using the exogenous variation in the geological stratum. In this approach, equation  2 is regarded as a structural form equation because the least-squares estimator ($\hat{\beta}$) is assumed to be attenuated by the measurement error in the exposure variable. The reduced-form equation for MineDeposit is designed as follows:
@@
where Stratum is a binary IV that equals one for municipalities with sedimentary rock created during the Cenozoic era, and $\epsilon$ is a random error term. The IV is plausibly excluded from the structural equation because the location of mines is essentially dominated by the distribution of geological stratum, which is exogenously given in nature and unobservable by people in the prewar period. In addition, the relevance condition holds because the location of coalfields is basically determined by the distribution of the stratum created in the Cenozoic era, which includes the Carboniferous period when the strata containing coal were created (Fernihough and O’Rourke 2020). Online Appendix C.2 summarizes the geological strata variable in finer detail and shows that the location of coal mines is determined by specific strata created in the Cenozoic era. Online Appendix D.2 discusses the validity of identification assumptions in more rigorous way.
I use the heteroskedasticity-consistent covariance matrix estimator as a baseline estimator. [^35] The results are materially similar between the Eicker-White type covariance matrix estimator suggested by Hinkley (1977) and the HC2 estimator proposed by Horn et al. (1975). For the reduced form regressions, to be conservative, I use the HC2 estimator given that the treatment variable is relatively sparse (Section 4.1). For the sensitivity check, I also use the standard errors clustered at the county level based on the cluster-robust covariance matrix estimator (Arellano 1987) to determine the influence of the potential influences of the local-scale spatial correlations. Online Appendix D.3 shows that the results are materially similar under both variance estimators. Therefore, the potential spatial correlations were negligible in this empirical setting.
@@ 5.2 Main Results
Notes: Panels A and B present the results for the 1930 and 1935 samples, respectively. Columns 1–3 show the estimates based on the IV estimator for the system of equations 2 and 3. Column 4 shows the estimates from the OLS estimator for equation 2. Panel A of Online Appendix Table C.4 shows the summary statistics for the population (before log-transformed) by treatment and control groups.
@@
Panel A of Table 2 presents the main results for the population in 1930. The estimate from a simple IV estimation is shown in column 1.363636The first-stage $F$-statistics reported in the same panel exceed the rule-of-thumb threshold value proposed by Staiger and Stock (1997), say 10, in all regressions; this shows that the necessary rank condition is satisfied. Online Appendix D.3 confirms that the first-stage $F$-statistics never break this threshold when I use a conservative variance estimator such as the cluster-robust variance-covariance matrix estimator. Then I include the city and town fixed effects (column 2) and both the fixed effects and control variables (column 3). The estimated coefficient decreases from $1.386$ to $0.829$ when I include the fixed effects. This suggests that, as expected, the location of the coal mines may be positively correlated with the local infrastructure. Including additional control variables on transportation accessibility has a moderate influence on the estimate ($0.682$). This supports the evidence that, after conditioning on the fixed effects, there are few influential unobservables that are potentially correlated with the location of mines. Panel B confirms the similar results for the population in 1935 in the same column layouts. The estimated coefficient from the specification including the full set of controls in column 3 is $0.723$, which is larger than that for the 1930 sample.
@@
The estimated magnitude is economically meaningful. The estimates from the preferred specification of column (3) indicate that coal mines increased the local population by $125$% [^37] Statistically, it is more precise to state that the geometric mean of population in the treatment group with mines was approximately 98% ($\text{exp}(0.682)=\text{exp}(\hat{\alpha}+0.682)/\text{exp}(\hat{\alpha})\approx 1.98$) greater than that of the control group. However, I use simple interpretations throughout this study to avoid redundancy. in 1930 (panel A) and $106$% in 1935 (panel B). The average number of miners per coal mine increased just less than $100$ during this period ($1,567$ in 1931 and $1,658$ in 1936). [^38] This is because while the number of male miners increased, while the number of female miners decreased due to the revision of the regulation (Table 1). In this light, it is a bit surprising that the magnitude is estimated to be greater in the 1935 sample. I will assess the mechanism behind this trend in the next subsections.
@@
In column 4, I present the estimates based on the reduced-form assumption to see if the IV approach works as expected. Even after conditioning on the control variables, the estimates based on the IV approach in column 3 are greater than those in column 4. This implies that my IV estimation strategy deals with systematic attenuation due to measurement errors.

### v14

@@ 4.3 Dependent Variables
The primary dependent variable, population, was obtained by digitizing the municipal-level statistics documented in the 1930 and 1935 Population Censuses (Statistics Bureau of the Cabinet 1935, 1939).
@@
To investigate the natural change mechanism behind the local population growth due to mining, the Municipal Vital Statistics of 1930 and 1935 (Statistics Bureau of the Cabinet 1932c, 1938) were digitized. They documented the total number of marriages, live births, and deaths in all municipalities in the survey years. Four dependent variables were then considered: crude marriage rate, crude birth rate, marital fertility rate, and gender-specific death rate. [^27] The conventional definitions of these variables were followed. The crude marriage and birth rates are the number of marriages and live births per $1,000$ people, respectively. Marital fertility is defined as the number of live births per $1,000$ married females. Although the number of married females in each age is unavailable, the number of married females is similar to the number of households in each municipality, meaning that the marital fertility definition herein is useful in representing the couple’s fertility decision. In other words, the results from this marital fertility definition shall not be sensitive to the existence of the elder married women in the households. Although marital fertility is considered to capture the couple’s fertility decision better, the results can be materially similar between crude and marital fertilities because out-of-wedlock births have been historically rare in Japan. The crude death rate is the number of deaths per $1,000$ people. These are used to test the fertility responses of the miners and the changes in mortality due to mining activities and regulations. To further analyze the occupational hazard channel, the infant mortality rate and fetal death rate were considered. [^28] The infant mortality rate is the number of infant deaths per $1,000$ live births. The number of censored observations in infant mortality is negligible. The fetal death rate is defined as the number of fetal deaths per $1,000$ births. The censored observations in the fetal death rates are less than $11$% in both census years. The estimates from the Tobit estimator (Tobin 1958) are materially similar to those from the OLSE (not reported). This confirms the evidence that censoring is not material in the empirical setting. These rates allow for testing the impacts of the occupational hazards for female miners on the mortality selection before birth under the biological framework (Section 5.4.1). The mortality rate of children aged 1–4 was also considered to assess the potential impacts of air pollution. [^29] The child mortality rate is defined as the number of deaths of children aged 1–4 in 1933 (or 1938) per $1,000$ children aged 1–5 in 1930. Note that the number of children aged 1–5 from 1930 is used as it is only documented in the 1930 population census. Given that child deaths, especially at age 5, were rarer than infant deaths at that time, the mismatch in the age bins between numerator and denominator should not be a critical problem. The difference in the survey points does add noise to the estimation of the standard error. However, the coefficient estimate is sufficiently close to zero and thus, should not be a practical issue in this setting (Table 5). To obtain both variables, the 1933 and 1938 reports documenting municipal-level statistics on births and infant deaths published by the Aiikukai and Social Welfare Bureau were digitized.
@@
Comprehensive statistics on migrants were unavailable in the prewar period. As described earlier, however, once the contribution of natural change to population growth is known, the sign of social change in the geometric population growth model is uniquely determined. Considering this, it is prefereable to assess the extent the influx of single male miners disturbed the gender balance, average household size, and how it changed after the regulation was enforced. The data on the sex ratio and average household size are obtained from the official reports of the 1930 and 1935 Population Censuses. [^30] The sex ratio is the number of males divided by the number of females. The average household size is the number of people in households divided by the number of households.
@@
Overall, these variables are expected to represent the characteristics of local demographics (Angrist 2002; Abramitzky et al. 2011). Importantly, the census documents the total number of people in all municipalities in the census years. Similarly, the vital statistics are corrected under a national registration system that covers all the births and deaths in a certain year. [^31] The potential imprecision of birth data in prewar Japan was mostly eliminated by the 1920s (Drixler 2016). This comprehensiveness helps avoid sample selection issues in statistical inference. Online Appendix Table C.4 presents the summary statistics on these variables by treatment status and census year. The mean differences are statistically significant for most variables, implying that systematic demographic changes may have occurred around the mining extraction area.
@@ 4.4 Control Variables
As coal is a bulky mineral, railways were the primary transportation of the extracted coal. Firms may have preferred setting extraction points closer to railway stations to reduce transportation costs (Sumiya 1968, pp. 440–441). Further, the local population size could be larger as one moves closer to stations. Therefore, controlling for the distance between each municipality and the nearest-neighboring station is preferable to deal with the potential endogeneity issue. The location information of all stations obtained from the official dataset provided by the Ministry of Land, Infrastructure, Transport and Tourism was used to compute the nearest-neighbor distance to stations in 1931 and 1936.
Although railways were primarily used to transport coal because coal mines were usually inland, marine transportation was also used for secondary logistics (Online Appendix Figure 2). Hence, similar to railway transportation, the distance to the nearest neighboring seaport is included as a variable to control for accessibility to marine transportation.
Accessibility to rivers is also considered. Not all mining areas have large rivers suitable for water transportation, meaning that rivers did not dominate the locations of coal mines. However, several mines used river transport as primary logistics until the development of railways before the late 1920s (Chiba and Yamada 1964). This implies that the spatial distribution of the river could have influenced the location of mines in such mining areas. To control for this possibility, the distance to the nearest neighboring river is included as a control variable.
Finally, the potential influence of the elevation is taken into account. The mine’s location might have been correlated with the elevation of municipalities because terrain may relate to the locational fundamentals such as the climatic conditions and fixed cost of transportation. The grid cell elevation data provided by the Ministry of Land, Infrastructure, Transport, and Tourism was used to systematically calculate the average elevation in each municipality.
As explained below, including these control variables has little influence on the estimates after conditioning on the city and town fixed effects. To be conservative, however, it is preferable to include all the controls to ensure the orthogonality condition of the instrumental variable. Online Appendices C.4.1C.4.4 summarizes the details of the data on railways, seaports, rivers, and elevation. Online Appendix Table C.5 presents the summary statistics on these control variables. [^32] Online Appendix Figures C.5C.6, and C.7 illustrate the locations of the railway stations, seaports, and rivers in the Japanese archipelago, respectively. Online Appendix Figure C.8 shows the spatial distribution of the elevation. All the controls are log-transformed to better model the skewness of the distances and elevations in nature, albeit this transformation does not influence the results.
@@ 5.1 Estimation Strategy
The random nature of mineral deposits is leveraged to identify the effects of mines. The linear regression model is as follows:
@@
where $i$ indexes municipalities, $y$ is the outcome variable, MineDeposit is an indicator variable that equals one for municipalities within 5 km from a mine, $\mathbf{x}$ is a vector of control variables, and $e$ is a random error term. As the placement of a mine is determined by a geological anomaly, MineDeposit is random in nature (Benshaul-Tolonen 2019, p.1568). However, the rest of the variation may be correlated with the local variation in infrastructure. Thus, if a mineral deposit is found between a village and a city, the mining firm may choose the city as its main mining point because the city is likely to have better infrastructure than the village. Then, MineDeposit can be positively correlated with the error term, leading to a positive (negative) omitted variable bias in the estimate of $\beta$ if the placement of the city is positively (negatively) correlated with the outcome variable.333333 As an example, consider a simplified projection of $y$ on MineDeposit and City, $y=\alpha+\beta\text{{MineDeposit}}+\gamma\textit{City}+e$. When the municipality type (City) is omitted, the linear projection coefficient can be written as $\beta^{*}=\beta+\Xi\gamma$, where $\Xi=(E[\text{{MineDeposit}}^{2}]^{-1})(E[\text{{MineDeposit}}\text{{City}})$. As explained, MineDeposit and City may be positively correlated such that $\Xi>0$. In addition, when $y$ is the population, conditional on mine deposits, the city has a greater number of people than town/villages ($\gamma>0$). This implies that $\beta^{*}=\beta+\Xi\gamma>\beta$. To deal with this systematic bias in the ordinary least squares, two indicator variables are included that equal one for cities and towns, respectively. As explained in Section 4.4, the distances to the nearest neighboring railway station, seaport, and river and the elevation are included to explicitly control for transportation accessibility. [^34] The differences in the estimates from both specifications, including and excluding the control variables, indicate how much this omitted variable mechanism influences the results. Thus, the randomness of the primary exposure variable can be partially assessed (MineDeposit). As demonstrated later, the results from the specification including the city and town fixed effects are materially similar to those from the specifications. This includes the control variables, thereby supporting the randomness of the exposure variable. To be conservative, the specification including the control variables is preferred.
@@
A potential threat in the identification may be measurement errors in time-dimensional assignments. As explained in Section 4.1, mine data were surveyed in October 1931 (1936), whereas the Population Census was conducted in October 1930 (1935). Generally, this lag concerning the censuses should not lead to a strong attenuation issue because the increments in the number of mining municipalities per year were considerably limited (Section 4.1). In turn, the data on fetal death and infant mortality were obtained from the 1933 (1938) Vital Statistics. The lags in the matching allow the consideration of the exposure durations. Thus, fetus miscarriages in 1933 (1938) were in utero conceived in 1932 (1937), whereas their mothers should have been exposed to any shocks at the time of conception in 1932 (1937). The same argument can be applied to infant mortality. Infants who died in 1933 (1938) were born at the beginning of 1932 (1937) at the earliest. Therefore, they should have been in utero in 1931 (1936). Consequently, the mining data that list the mines in October 1931 and 1936 may be reasonably matched with the health data of 1933 and 1938, respectively. However, one must be careful as there is still a one-year lag in matching with both the census and vital statistics datasets. This may lead to attenuation bias due to miss assignments in the time dimension, because the number of coal mines should be increased during that year. [^35] Note that this type of measurement error never overstates the estimates but causes attenuation. Thus, if a few treated (untreated) municipalities were regarded as the untreated (treated) municipalities, the impacts of mines shall be discounted. See Online Appendix D.1 for a brief explanation of this mechanism.
@@
To deal with such potential issues, the IV estimator is considered using the exogenous variation in the geological stratum. In this approach, equation  2 is regarded as a structural form equation because the least-squares estimator ($\hat{\beta}$) is assumed to be attenuated by the measurement error in the exposure variable. The reduced-form equation for MineDeposit is designed as follows:
@@
where Stratum is a binary IV that equals one for municipalities with sedimentary rock created during the Cenozoic era, and $\epsilon$ is a random error term. The IV is plausibly excluded from the structural equation because the location of mines is essentially dominated by the distribution of geological stratum, which is exogenously given in nature and unobservable by people in the prewar period. In addition, the relevance condition holds because the location of coalfields is basically determined by the distribution of the stratum created in the Cenozoic era, which includes the Carboniferous period when the strata containing coal were created (Tanaka 1984; Fernihough and O’Rourke 2020). Online Appendix C.2 summarizes the geological strata variable in finer detail and shows that the location of coal mines is determined by specific strata created in this era. Online Appendix D.2 discusses the validity of identification assumptions in more rigorous way.
The heteroskedasticity-consistent covariance matrix estimator is used as a baseline estimator. [^36] The results are materially similar between the Eicker-White type covariance matrix estimator suggested by Hinkley (1977) and the HC2 estimator proposed by Horn et al. (1975). For the reduced form regressions, to be conservative, the HC2 estimator is used given that the treatment variable is relatively sparse (Section 4.1). For the sensitivity check, the standard errors clustered at the county level based on the cluster-robust covariance matrix estimator (Arellano 1987) are used to determine the influence of the potential influences of the local-scale spatial correlations. Online Appendix D.3 shows that the results are materially similar under both variance estimators. Therefore, the potential spatial correlations were negligible in this empirical setting.
@@ 5.2 Main Results
Notes: Panels A and B present the results for the 1930 and 1935 samples, respectively. Columns 1–3 show the estimates based on the IV estimator for the system of equations 2 and 3. Column 4 shows the estimates from the OLS estimator for equation 2. Geometric Mean in the TG and CG show the $\text{exp}(\hat{\alpha}+\hat{\beta})$ and $\text{exp}(\hat{\alpha})$ from the estimated system, respectively. Panel A of Online Appendix Table C.4 shows the summary statistics for the population (before log-transformed) by treatment and control groups.
@@
Panel A of Table 2 presents the main results for the population in 1930. The estimate from a simple IV estimation is shown in column 1.373737The first-stage $F$-statistics reported in the same panel exceed the rule-of-thumb threshold value proposed by Staiger and Stock (1997), say 10, in all regressions. This shows that the necessary rank condition is satisfied. Online Appendix D.3 confirms that the first-stage $F$-statistics never break this threshold when a conservative variance estimator such as the cluster-robust variance-covariance matrix estimator is used. Then I include the city and town fixed effects (column 2), and both the fixed effects and control variables (column 3). The estimated coefficient decreases from $1.386$ to $0.829$ when the fixed effects are included. This suggests that, as expected, the location of the coal mines may be positively correlated with the local infrastructure. Including additional control variables on transportation accessibility has a moderate influence on the estimate ($0.682$). This supports the evidence that, after conditioning on the fixed effects, there are few influential unobservables that are potentially correlated with the location of mines. Panel B confirms the similar results for the population in 1935 in the same column layouts. The estimated coefficient from the specification including the full set of controls in column 3 is $0.723$, which is larger than that for the 1930 sample.
@@
The estimated magnitude is economically meaningful. The estimates from the preferred specification of column (3) indicate that coal mines increased the local population by $98$% [^38] Statistically, it is more precise to state that the geometric mean of population in the treatment group with mines was approximately 98% (i.e., $\text{exp}(0.682)=\text{exp}(\hat{\alpha}+0.682)/\text{exp}(\hat{\alpha})\approx 1.98$) greater than that of the control group. However, simple interpretations are used throughout this study to avoid redundancy. in 1930 (panel A) and $106$% in 1935 (panel B). The average number of miners per coal mine increased just less than $100$ during this period ($1,567$ in 1931 and $1,658$ in 1936). [^39] This is because while the number of male miners increased, simultaneously the number of female miners decreased due to the revision of the regulation (Table 1). From this perspective, it is surprising that the magnitude is estimated to be greater in the 1935 sample. The mechanism behind this trend is assessed in the next subsections.
@@
Column 4 presents the estimates based on the reduced-form assumption to determine whether the IV approach works as expected. Even after conditioning on the control variables, the estimates based on the IV approach in column 3 are greater than those in column 4. This implies that the IV estimation strategy deals with systematic attenuation due to measurement errors.